# AI/ML based Joint Source and Channel Coding for HARQ-ACK Payload

### v1

@@ VI-C Low Complexity Coherent Receiver
Both the receivers proposed in Section VI-A and VI-B compute identical probabilities. The coherent design is more compatible with traditional receiver processing whereby we first use the pilot REs to perform channel estimation (ChEST) and then utilize the channel estimate on the data REs to decode (provided $\beta_{m}=1$). However, the computation of the bitwise likelihood in (19) using the receivers described thus far will require computing a weighted sum of $2^{k-1}$ exponentials for each of the $k$ information bits. In order to reduce the computational complexity, we demonstrate how to simplify the coherent receiver design under the following assumptions
  • $p(\mathbf{y}_{p,l}|\mathcal{H}_{m})$ can be dropped i.e. $\beta_{m}=1$
  • $\mathbf{R}_{\mathbf{hh}}=\mathbf{1}_{N}\mathbf{1}_{N}^{T}$ i.e., channel is frequency flat.
Our results in Section IX-D will demonstrate that we can drop $p(\mathbf{y}_{p,l}|\mathcal{H}_{m})$ even if $\beta_{m}\neq 1$ without significant performance degradation. However, the second assumption will cause minimal performance degradation only if the number of resource blocks (RBs) allocated is small so that the channel is effectively “frequency-flat”. This is typically the case for PUCCH transmissions.
@@
is the Mahalanobis distance between $\mathbf{y}_{d,l}$ and $\boldsymbol{\mu}_{d,l}$. By substituting for flat fading in $\mathbf{R}_{\mathbf{hh}}$ and repeatedly utilizing the matrix inversion lemma (1) (refer Appendix E for the proof), we obtain
$\begin{gathered}\mathrm{det}\mathbf{\Sigma}_{d,l}=\mathrm{NV}\bigg(1+\frac{\alpha_{m}N_{d}}{\beta_{m}N_{p}+\frac{1}{\mathrm{SNR}}}\bigg)\\ \Delta^{2}(\mathbf{y}_{d,l},\boldsymbol{\mu}_{d,l})=\frac{1}{\mathrm{NV}}\bigg(\|\mathbf{y}_{d,l}-\boldsymbol{\mu}_{d,l}\|^{2}_{2}-\frac{\alpha_{m}}{\alpha_{m}N_{d}+\beta_{m}N_{p}+\frac{1}{\mathrm{SNR}}}\|(\mathbf{y}_{d,l}-\boldsymbol{\mu}_{d,l})^{H}\mathbf{c}_{m}\|^{2}_{2}\bigg).\end{gathered}$ (36)
Observe that the definition of $\Delta^{2}(\mathbf{y}_{d,l},\boldsymbol{\mu}_{d,l})$ in (36) contains both $2\mathrm{Re}(\mathbf{y}_{d,l}^{H}\boldsymbol{\mu}_{d,l})$ (from $\|\mathbf{y}_{d,l}-\boldsymbol{\mu}_{d,l}\|^{2}_{2}$) and $\|\mathbf{y}_{d,l}^{H}\mathbf{c}_{m}\|^{2}_{2}$ (from $\|(\mathbf{y}_{d,l}-\boldsymbol{\mu}_{d,l})^{H}\mathbf{c}_{m}\|^{2}_{2}$). Recalling that $\boldsymbol{\mu}_{d,l}$ is an elementwise product of the channel estimate and the codeword $\mathbf{c}_{m}$, the first term models traditional coherent communication e.g. codeword ML decoder. Meanwhile, the second term computes the energy of $\mathbf{y}_{d,l}$ when projected onto $\mathbf{c}_{m}$, which is typically computed by an energy detector used for non-coherent communication.
@@
Computational Complexity Analysis: Returning to (34), we observe that $\log{\frac{\pi_{m}}{\mathrm{det}\mathbf{\Sigma}_{d,l}}}$ can be precomputed $\forall~m\in[2^{k}]$. More precisely, $\log{\frac{\pi_{m}}{\mathrm{det}\mathbf{\Sigma}_{d,l}}}+\log{\mathrm{NV}}$ is independent of the received signal (since we will fix SNR used in receiver processing to a nominal value independent of the received SNR as will be detailed in Section IX-B) and adding $\log{\mathrm{NV}}$ is $\mathcal{O}(1)$. Since $\mathbf{y}_{d,l}$, $\boldsymbol{\mu}_{d,l}$ and $\mathbf{c}_{m}$ are vectors of length $n/2$, computing $\|\mathbf{y}_{d,l}-\boldsymbol{\mu}_{d,l}\|^{2}_{2}$ and $\|(\mathbf{y}_{d,l}-\boldsymbol{\mu}_{d,l})^{H}\mathbf{c}_{m}\|^{2}_{2}$ are both $\mathcal{O}(n)$. Furthermore, we compute $\Delta^{2}(\mathbf{y}_{d,l},\boldsymbol{\mu}_{d,l})$ for each of the $N_{r}$ receive antennas and $2^{k}$ codeword hypothesis – $\mathcal{O}(2^{k}N_{r}n)$. Finally, to evaluate (37), we have to compute a max over the $2^{k-1}$ values of $\Delta^{2}(\mathbf{y}_{d,l},\boldsymbol{\mu}_{d,l})+\log{\frac{\pi_{m}}{\mathrm{det}\mathbf{\Sigma}_{d,l}}}$ for $b_{i}=0$ and $b_{i}=1$, and repeat this for each of the $k$ information bits – $\mathcal{O}(k2^{k})$. Hence, the computational complexity is $\mathcal{O}(2^{k}(N_{r}n+k))$, compared to an ML decoder which is $\mathcal{O}(2^{k}N_{r}n)$. Hence for $k\ll nN_{r}$, the legacy ML decoder and the low complexity coherent receiver have the same order of computational complexity.
@@ VII-A Probabilistic Model for HARQ-ACK distribution
While it can be argued that a first order Markov model may not be sufficient to capture the correlation among HARQ-ACK bits, PDSCH CRC logs from field measurements have supported the adequacy of first order Markov modelling. However, we would like to emphasise that our encoder and decoder design principles can be applied to any higher order Markov model or more complex probability distributions, as long as the HARQ-ACK bits are sufficiently biased towards ACK.
Figure 4: Transition probabilites for consecutive HARQ-ACK bits
@@ VII-D Loss Functions & Training Algorithms
As part of the transformer training, the transformer will encode a batch of message sequences which are subsequently power shaped, passed through a SISO AWGN channel and decoded assuming perfect channel estimation. Since we are focused on the unequal protection of ACK and NACK bit errors as introduced in Section V, our training obective should attempt to lower both the ACK and NACK bit error rate. Let us first focus on reducing $\mathrm{BER}_{\mathrm{ACK}}$ at a given SNR. Observe that
@@ VIII AWGN Receiver Results
VIII-A Computing $\mathrm{SNR}_{\mathrm{th}}$
@@
VIII-B Effect of varying $\rho$ on $\mathrm{SNR}_{\mathrm{th}}$
@@
VIII-C Comparing SS vs FL training
@@ IX Fading Channel Receiver Results
Taking into account the practical considerations on the encoder side, in accordance with the results presented in Section VIII, we will utilize a quantized transformer codebook for $(k,n)=(4,32)$ with Step PS of 3 dB and a Bitwise UEP receiver777k=4k=4 was determined to be the most common value for HARQ-ACK in external discussions with other companies. . The transformer utilizes $A=2$ tokens of input length 2 each. We use a relatively low value of Step PS to ensure a low power jump across OFDM symbols, if the DMRS is not power shaped and the DMRS and data are time-division multiplexed on consecutive OFDM symbols. A Bitwise UEP receiver is utilized in place of the more practical MLM GLRT receiver, with the understanding that MLM GLRT will lead to 0.30.40.3-0.4 dB degradation for all receivers (excluding the ML baseline). This is because imprecise computations of the codeword likelihoods $p(\mathbf{y}|\mathcal{H}_{m})$ will require the utilization of different $\beta$ values for different algorithms, which could be hard to analyze with MLM GLRT also being utilized.
@@ IX-B Computing receive SNR and noise variance in practice
While we can assume genie knowledge of $\mathrm{SNR}$ and $\mathrm{NV}$ in simulation, these will have to be computed from reference signals in practice. For computing the SNR, we have two options: it can be estimated from the DMRS REs, or we can assume a nominal value. Given only 8 DMRS REs, the SNR estimated from the DMRS may have high variance. Hence we assume a fixed nominal SNR at the receiver (irrespective of the actual received SNR). In Appendix F, we will present empirical evidence that demonstrates minimal degradation from overestimation of the received SNR. Since PUCCH performance gains are to be obtained at low SNR i.e. <0<0 dB, we will fix the nominal SNR to $\mathrm{SNR}_{\mathrm{n}}=0$ dB. Given that the DMRS is shaped, for hypothesis $\mathcal{H}_{m}$, the hypothesized nominal SNR is $\beta_{m}\mathrm{SNR}_{\mathrm{n}}$.
Estimating $\mathrm{NV}$ can be performed subsequent to channel estimation. Per Rx antenna, we can estimate $\mathbb{E}[\mathbf{n}_{l}^{2}]=\|\mathbf{y}_{p,l}-\mathbf{\hat{h}}_{p,l}\odot\mathbf{c}_{p,l}\|^{2}_{2}/N_{p}$, and average $\mathbb{E}[\mathbf{n}_{l}^{2}]$ over all receive antennas to obtain an estimate of $\mathrm{NV}$.
@@ IX-C Channel Estimation
Since we are evaluating over 1 RB, we can also consider the channel to be approximately frequency flat and set all entries of $\mathbf{R}_{\mathbf{hh}}$ to be 1 while performing ChEST with negligible performance degradation.
@@ IX-D Comparing coherent and noncoherent receiver designs
In Fig. 12a, we plot $\mathrm{SNR}_{\mathrm{th}}$ vs $\rho$ for various schemes under a fading channel (TDL-C 300 ns). The NR + ML baseline is a correlator based decoder i.e. we compute $\mathrm{Re}(\sum_{l}\mathbf{y}^{H}_{d,l}(\mathbf{\hat{h}}_{d,l}\odot\mathbf{c}_{m}))$ for each hypothesized codeword $\mathbf{c}_{m}$ and find the codeword that maximizes the correlation. Since the ML decoder is unimpacted by $\rho$, $\mathrm{SNR}_{\mathrm{th}}$ is around 0.2 dB for all $\rho$. This is the coherent ML decoder that would be utilized in practice. However, the optimum codeword ML decoder would be noncoherent and involve computing $p(\mathbf{y}_{l}|\mathcal{H}_{m})$ using (20) and then finding the codeword index
$m^{*}=\arg\max_{m}p(\mathbf{Y}|\mathcal{H}_{m}).$ (47)
This improves $\mathrm{SNR}_{\mathrm{th}}$ by 2 dB. If we utilize the Bitwise UEP receiver in place of the codeword ML decoder, having computed $p(\mathbf{y}_{l}|\mathcal{H}_{m})$ optimally, a further 1 dB improvement in $\mathrm{SNR}_{\mathrm{th}}$ is obtained. Finally, a 1 dB improvement can be obtained from Step PS with $\delta=3$ dB and 0.50.70.5-0.7 dB gain can be obtained from employing the TF-based encoder. Note that $\delta=3$ dB is a very pessimistic value, and further gains can be obtained from PS by increasing $\delta$, while ensuring $P_{\mathrm{max}}$ is less than $\mathrm{SNR}_{\mathrm{th}}$ of the NR + ML baseline. Since the optimal coherent and noncoherent decoder are theoretically identical, we confirm via simulation that the $\mathrm{SNR}_{\mathrm{th}}$ obtained from both line up. Replacing the optimal coherent decoder by its low complexity approximation from Section VI-C leads to a minimal degradation of $0.3$ dB for both TF and NR.
@@ X Conclusions and Future Directions
Since NR channel codes have been designed assuming a uniform source prior, they are unable to exploit the biased HARQ-ACK source prior which arises from the PDSCH BLER being maintained under $10\%$. In this paper, we make fundamental contributions to joint source and channel coding (and modulation) for HARQ-ACK bits transmitted in the uplink. On the encoder side, we propose a non-linear transformer-based codebook to exploit the source prior. Furthermore, we demonstrate how assigning less power to more frequently seen codewords and vice-versa can lower the SNR required to achieve the target error rates. On the decoder side, we propose the first extension of the Neyman Pearson test to a coded bit system with multiple information bits to preferentially protect NACK over ACK. This further lowers the SNR needed to achieve a NACK and ACK error rate of 0.1% and 1% respectively. For both the encoder and decoder, we propose methods to lower complexity and improve robustness to small changes and/or partial knowledge of the HARQ-ACK distribution. Finally, we present comprehensive end-to-end NR-compliant simulations in an uplink SIMO setup and develop the receiver from first principles to demonstrate the effectiveness of JSCC for HARQ-ACK. By obtaining 3 – 6 dB reduction in average transmit power and 2 – 3 dB reduction in the maximum transmit power over the NR baseline while accounting for practical constraints, our approach provides for significant power savings and coverage gains.
As part of future work, a single codebook can be designed that works for a range of $(k,n)$ values, akin to the generator matrix used in [3gpp.38.212]. While we presented results for two $k$ values, $k=4$ and $k=11$, and designed a separate transformer for each, having a single codebook for $3\leq k\leq 11$ would be desirable. In this paper, we restrict ourselves to JSCC for HARQ-ACK. However, one could extend this approach to exploit temporal correlation in downlink control information. Furthermore, for larger payloads $k>11$, an ML/MAP decoder would become computationally infeasible, hence an AI based encoder would have to be coupled with an AI based decoder to ensure manageable computational complexity. We hope that the first principles based approach taken in this paper can inspire further innovation in the development of practical AI/ML based enhancements to the PHY layer for 6G and beyond.

### v2

@@ VI-C Low Complexity Coherent Receiver
Both the receivers proposed in Section VI-A and VI-B compute identical probabilities. However, the computation of the bitwise likelihood in (19) using the receivers described thus far will require computing a weighted sum of $2^{k-1}$ exponentials for each of the $k$ information bits. In order to reduce the computational complexity, we demonstrate how to simplify the coherent receiver design described in Section VI-B assuming $\beta_{m}=1$ and $\mathbf{R}_{\mathbf{hh}}=\mathbf{1}_{N}\mathbf{1}_{N}^{T}$ i.e., channel is frequency flat. Our results in Section IX-D will demonstrate that we can drop $p(\mathbf{y}_{p,l}|\mathcal{H}_{m})$ even if $\beta_{m}\neq 1$ without significant performance degradation. However, the second assumption will cause minimal performance degradation only if the number of resource blocks (RBs) allocated is small as is the case for PUCCH, so that the channel is effectively “frequency-flat”.
@@
is the Mahalanobis distance between $\mathbf{y}_{d,l}$ and $\boldsymbol{\mu}_{d,l}$. By substituting for flat fading in $\mathbf{R}_{\mathbf{hh}}$ and repeatedly utilizing the matrix inversion lemma (1) (refer Appendix F for the proof), we obtain
$\begin{gathered}\mathrm{det}\mathbf{\Sigma}_{d,l}=\mathrm{NV}^{N_{d}}\bigg(1+\frac{\alpha_{m}N_{d}}{\beta_{m}N_{p}+\frac{1}{\mathrm{SNR}}}\bigg)\\ \Delta^{2}(\mathbf{y}_{d,l},\boldsymbol{\mu}_{d,l})=\frac{1}{\mathrm{NV}}\bigg(\|\mathbf{y}_{d,l}-\boldsymbol{\mu}_{d,l}\|^{2}_{2}-\frac{\alpha_{m}}{\alpha_{m}N_{d}+\beta_{m}N_{p}+\frac{1}{\mathrm{SNR}}}\|(\mathbf{y}_{d,l}-\boldsymbol{\mu}_{d,l})^{H}\mathbf{c}_{m}\|^{2}_{2}\bigg).\end{gathered}$ (36)
Observe that the definition of $\Delta^{2}(\mathbf{y}_{d,l},\boldsymbol{\mu}_{d,l})$ in (36) contains both $2\Re(\mathbf{y}_{d,l}^{H}\boldsymbol{\mu}_{d,l})$ (from $\|\mathbf{y}_{d,l}-\boldsymbol{\mu}_{d,l}\|^{2}_{2}$) and $\|\mathbf{y}_{d,l}^{H}\mathbf{c}_{m}\|^{2}_{2}$ (from $\|(\mathbf{y}_{d,l}-\boldsymbol{\mu}_{d,l})^{H}\mathbf{c}_{m}\|^{2}_{2}$). Recalling that $\boldsymbol{\mu}_{d,l}$ is an elementwise product of the channel estimate and the codeword $\mathbf{c}_{m}$, the first term models traditional coherent communication e.g. codeword ML decoder. Meanwhile, the second term computes the energy of $\mathbf{y}_{d,l}$ when projected onto $\mathbf{c}_{m}$, which is typically computed by an energy detector used for non-coherent communication.
@@
Computational Complexity Analysis: Returning to (34), we observe that $\log{\frac{\pi_{m}}{\mathrm{det}\mathbf{\Sigma}_{d,l}}}$ can be precomputed $\forall~m\in[2^{k}]$. More precisely, $\log{\frac{\pi_{m}}{\mathrm{det}\mathbf{\Sigma}_{d,l}}}+\log{\mathrm{NV}}$ is independent of the received signal (since we will fix SNR used in receiver processing to a nominal value independent of the received SNR as will be detailed in Section IX-B). Given $\mathbf{y}_{d,l}$, $\boldsymbol{\mu}_{d,l}$ and $\mathbf{c}_{m}$ are vectors of length $n/2$, computing $\Delta^{2}(\mathbf{y}_{d,l},\boldsymbol{\mu}_{d,l})$ requires $9n/2+2$ MAC operations. Furthermore, we compute $\Delta^{2}(\mathbf{y}_{d,l},\boldsymbol{\mu}_{d,l})$ for each of the $N_{r}$ receive antennas and $2^{k}$ codeword hypothesis – $2^{k}N_{r}(9n/2+2)$. Finally, to evaluate (37), we have to compute a max over the $2^{k-1}$ values of $\Delta^{2}(\mathbf{y}_{d,l},\boldsymbol{\mu}_{d,l})+\log{\frac{\pi_{m}}{\mathrm{det}\mathbf{\Sigma}_{d,l}}}$ for $b_{i}=0$ and $b_{i}=1$, and repeat this for each of the $k$ information bits – $k2^{k}$ operations. Hence, the computational complexity is $2^{k}(N_{r}(9n/2+2)+k))$ MAC operations. An ML decoder computes $\sum_{l}\mathbf{y}^{H}_{d,l}(\mathbf{\hat{h}}_{d,l}\odot\mathbf{c}_{m})$ for each codeword $\mathbf{c}_{m}$, which requires $2^{k}(2nN_{r}+1)$ MAC operations. Hence for $k\ll nN_{r}$, the legacy ML decoder and the low complexity coherent receiver have comparable computational complexities.
@@ VII-A Probabilistic Model for HARQ-ACK distribution
While it can be argued that a first order Markov model may not be sufficient to capture the correlation among HARQ-ACK bits, PDSCH CRC logs from field measurements have supported the adequacy of first order Markov modeling.
Figure 4: Transition probabilities for consecutive HARQ-ACK bits
@@ VII-D Loss Functions & Training Algorithms
As part of the transformer training, the transformer will encode a batch of message sequences which are subsequently power shaped, passed through a SISO AWGN channel and decoded assuming perfect channel estimation. Since we are focused on the unequal protection of ACK and NACK bit errors as introduced in Section V, our training objective should attempt to lower both the ACK and NACK bit error rate. Let us first focus on reducing $\mathrm{BER}_{\mathrm{ACK}}$ at a given SNR. Observe that
@@ VIII AWGN Receiver Results
VIII-A Computing threshold SNR $\mathrm{SNR}_{\mathrm{th}}$
@@
VIII-B Effect of varying bit correlation $\rho$ on $\mathrm{SNR}_{\mathrm{th}}$
@@
VIII-C Comparing Scenario Specific (SS) vs Free Lunch (FL) training
@@ IX Fading Channel Receiver Results
Taking into account the practical considerations on the encoder side, in accordance with the results presented in Section VIII, we will utilize a quantized transformer codebook for $(k,n)=(4,32)$ with Step PS of 3 dB and a Bitwise UEP receiver101010k=4k=4 was determined to be the most common value for HARQ-ACK in external discussions with other companies. . The transformer utilizes $A=2$ tokens of input length 2 each. We use a relatively low value of Step PS to ensure a low power jump across OFDM symbols, if the DMRS is not power shaped and the DMRS and data are time-division multiplexed on consecutive OFDM symbols. A Bitwise UEP receiver is utilized in place of the more practical MLM GLRT receiver, with the understanding that MLM GLRT will lead to 0.30.40.3-0.4 dB degradation for all receivers (excluding the ML baseline). This is because imprecise computations of the codeword likelihoods $p(\mathbf{y}|\mathcal{H}_{m})$ will require the utilization of different $\beta$ values for different algorithms, which could be hard to analyze with MLM GLRT also being utilized.
@@ IX-B Computing receive SNR and noise variance in practice
While we can assume genie knowledge of $\mathrm{SNR}$ and $\mathrm{NV}$ in simulation, these will have to be computed from reference signals in practice. For computing the SNR, we have two options: it can be estimated from the DMRS REs, or we can assume a nominal value. Given only 8 DMRS REs, the SNR estimated from the DMRS may have high variance. Hence we assume a fixed nominal SNR at the receiver (irrespective of the actual received SNR). In Appendix G, we will present empirical evidence that demonstrates minimal degradation from overestimation of the received SNR. Since PUCCH performance gains are to be obtained at low SNR i.e. <0<0 dB, we will fix the nominal SNR to $\mathrm{SNR}_{\mathrm{n}}=0$ dB. Given that the DMRS is shaped, for hypothesis $\mathcal{H}_{m}$, the hypothesized nominal SNR is $\beta_{m}\mathrm{SNR}_{\mathrm{n}}$. Furthermore, per receive antenna, we can estimate $\mathbb{E}[\mathbf{n}_{l}^{2}]$ as $\|\mathbf{y}_{p,l}-\mathbf{\hat{h}}_{p,l}\odot\mathbf{c}_{p,l}\|^{2}_{2}/N_{p}$, and then average $\mathbb{E}[\mathbf{n}_{l}^{2}]$ over all receive antennas to obtain $\mathrm{NV}$.
@@ IX-C Channel Estimation
Since we are evaluating over 1 RB, we can also consider the channel to be approximately frequency flat and set all entries of $\mathbf{R}_{\mathbf{hh}}$ to be 1 while performing ChEST.
@@ IX-D Comparing coherent and noncoherent receiver designs
In Fig. 12a, we plot $\mathrm{SNR}_{\mathrm{th}}$ vs $\rho$ for various schemes under a fading channel (TDL-C 300 ns). The NR + ML baseline is a correlator based decoder i.e. we compute $\Re(\sum_{l}\mathbf{y}^{H}_{d,l}(\mathbf{\hat{h}}_{d,l}\odot\mathbf{c}_{m}))$ for each hypothesized codeword $\mathbf{c}_{m}$ and find the codeword that maximizes the correlation. Since the ML decoder is unimpacted by $\rho$, $\mathrm{SNR}_{\mathrm{th}}$ is around 0.2 dB for all $\rho$. This is the coherent ML decoder that would be utilized in practice. However, the optimum codeword ML decoder would be noncoherent and involve computing $p(\mathbf{y}_{l}|\mathcal{H}_{m})$ using (20) and then finding the codeword index $m^{*}=\arg\max_{m}p(\mathbf{Y}|\mathcal{H}_{m})$. This improves $\mathrm{SNR}_{\mathrm{th}}$ by 2 dB. If we utilize the Bitwise UEP receiver in place of the codeword ML decoder, having computed $p(\mathbf{y}_{l}|\mathcal{H}_{m})$ optimally, a further 1 dB improvement in $\mathrm{SNR}_{\mathrm{th}}$ is obtained. Finally, a 1 dB improvement can be obtained from Step PS with $\delta=3$ dB and 0.50.70.5-0.7 dB gain can be obtained from employing the TF-based encoder. Note that $\delta=3$ dB is a very pessimistic value, and further gains can be obtained from PS by increasing $\delta$, while ensuring $P_{\mathrm{max}}$ is less than $\mathrm{SNR}_{\mathrm{th}}$ of the NR + ML baseline. Since the optimal coherent and noncoherent decoder are theoretically identical [^11] As derived in Section VI-B, if we utilize an LMMSE channel estimator whilst utilizing the optimal coherent decoder, then $p(\mathbf{y}_{l}|\mathcal{H}_{m})$ computed in Section VI-B and VI-A are identical., we confirm via simulation that the $\mathrm{SNR}_{\mathrm{th}}$ obtained from both line up. Replacing the optimal coherent decoder by its low complexity approximation leads to a minimal degradation of $0.3$ dB for TF and NR.
@@ X Conclusions and Future Directions
Since NR channel codes have been designed assuming a uniform source prior, they do not jointly perform source encoding that could exploit the biased HARQ-ACK source prior which arises from the PDSCH BLER being maintained under $10\%$ [^12] Even with the current NR codebook, the decoder could always utilize the HARQ-ACK prior to provide power savings.. In this paper, we make fundamental contributions to joint source and channel coding (and modulation) for HARQ-ACK bits transmitted in the uplink. On the encoder side, we propose a non-linear transformer-based codebook to exploit the source prior. Furthermore, we demonstrate how assigning less power to more frequently seen codewords and vice-versa can lower the SNR required to achieve the target error rates. On the decoder side, we propose the first extension of the Neyman Pearson test to a coded bit system with multiple information bits to preferentially protect NACK over ACK. This further lowers the SNR needed to achieve a NACK and ACK error rate of 0.1% and 1% respectively. For both the encoder and decoder, we propose methods to lower complexity and improve robustness to small changes and/or partial knowledge of the HARQ-ACK distribution. Finally, we present comprehensive end-to-end NR-compliant simulations in an uplink SIMO setup and develop the receiver from first principles to demonstrate the effectiveness of JSCC for HARQ-ACK. By obtaining 3 – 6 dB reduction in average transmit power and 2 – 3 dB reduction in the maximum transmit power over the NR baseline while accounting for practical constraints, our approach provides for significant power savings and coverage gains.
As part of future work, a single codebook can be designed that works for a range of $(k,n)$ values, akin to the generator matrix used in [3gpp.38.212]. While we presented results for two $k$ values, $k=4$ and $k=11$, and designed a separate transformer for each, having a single codebook for $3\leq k\leq 11$ would be desirable. For larger payloads $k>11$, an ML/MAP decoder would become computationally infeasible, hence an AI based encoder would have to be coupled with an AI based decoder to ensure manageable computational complexity. On a theoretical front, proving the optimality of $\beta$ as defined in (15) remains an open question. We hope that the first principles based approach taken in this paper can inspire further innovation in the development of practical AI/ML based enhancements to the PHY layer for 6G and beyond.